OpenStudy (anonymous):

-2 2/3 divided by 3/4

OpenStudy (anonymous):

dont understand

OpenStudy (anonymous):

Can you rewrite -2 2/3?

OpenStudy (mertsj):

He means change it to an improper fraction

OpenStudy (mertsj):

Remember dividing is the same as multiplying by the reciprocal.

OpenStudy (anonymous):

ya i know that but after that i dont know hat to do

OpenStudy (anonymous):

(a/b) / (c/d) = (a/b) * (d/c)

OpenStudy (anonymous):

- 2 2/3 as an improper fration is -8/3 now we have -8/3 divided by 3/4

OpenStudy (anonymous):

ok thanks everyone :)

OpenStudy (anonymous):

= -8/3 ÷ 3/4 = (-8 × 4) / (3 × 3) = -32/9 = -32/9 = -3 5/9

OpenStudy (anonymous):

there your answer

OpenStudy (anonymous):

thanks